Dear Richard dear all,

please find below a short summary of the activities today on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: A. Puglisi, P. Stefanini, E. Pinna, A. Riccardi, R. Briguglio, S. Esposito, L. Busoni, M. Xompero
LBTO: G. Brusa, and telescope crew.
MG: M. Andrighettoni
ADS: E. Anaclerio

Total # of person: 11

W unit

1) WFS cover. The cover has been installed and some measurements have been taken on the fabric. The cover will be tailored by B. Ferguson. The upgraded cover should be ready by tomorrow.
2) The IRTC has been bolted to the AGW unit. The two units now are standing on the AGW chart while the lab crane is supporting the IRTC.
3) some activity on WFS board alignment with teh AGW flange has been done. Data reduction is in progress.

AO control SW

1) Pyro module for remote interferometer control has been installed and tested on both the Interferometer PC and the AdSec control computer.


LBT672a

1) Installation of reference fiber for pupil alignment. The fiber are switched on and off via ethernet.


Next update tomorrow

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback